Week of 180305

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Kate(DB, chair), Julia (WLCG), Maarten (ALICE, WLCG), Remy (storage), Vincent (sec), Alberto (mon), Borja (mon), Gavin (comp), Marian (network)

  • remote: Alexander (NL-T1), Dmytro (NDGF), Kyle (OSG), Marcelo (CNAF), Sang Un (KISTI), Stephan (CMS), Di Qing (TRIUMF), Darren (RAL), David (IN2P3), Pepe (PIC), Victor (JINR), Xavier (KIT), Zoltan (LHCb)

Experiments round table:

  • CMS reports ( raw view) -
    • No major grid computing issue(s)
    • Running smoothly at a little over 200k cores
      • 60% production, 40% analysis
    • Tier-0 migration, OpenStack to HTCondor-CEs) in progress
    • Singularity deployment progressing well
      • large ticket for factory operations to correct entries (glexec/OS) for CEs where sites did not contact them after Singularity setup
    • tickets for all Tier-1,2 sites without Singularity and no known plan/schedule
    • CMS global cosmic run starting today
    • may not be able to attend the meeting but will read minutes, sorry!

  • ALICE -
    • Lowish activity level on average
      • New productions are being prepared

  • LHCb reports ( raw view) -
    • Activity
      • HLT farm fully running
      • 2017 data re-stripping ongoing
      • Stripping 29 reprocessing is ongoing
    • Site Issues
    • Tier2D
      • Users with UK certificate are having problem to access data at CBPF, Glasgow, CSCS, NCBJ (3 DPM, 1 dCache) GGUS:133667, GGUS:133617
Maarten offered to check the certificate issue.

Sites / Services round table:

  • ASGC: nc
  • BNL: nc
  • CNAF:
    • LHCb data move from old SE to new SE in progress, should finish by middle~end of this week. Also renewal of machine certificates.
    • Shared file system works in progress, the farm should be re-opened today or tomorrow.
    • Singularity is installed in CINECA farm.

Service VO Status Expected restart date Readiness GGUS ticket CNAF comment VO comment
Electric power line - Maint. 20.02, 21.02 One line in Production with UPS   The second redundant line is still down  
Tape buffer ALICE OK   Production 133582 all ALICE services at CNAF are running in the final configuration. Looks OK
Tape buffer ATLAS OK   Production 131742 Ok  
Tape buffer CMS OK   Production 133515 Ok  
Tape buffer LHCb Being Rebuilt Production in March Not Ready 133673    
Disk ALICE Parity OK   Production 133582 all ALICE services at CNAF are running in the final configuration. Looks OK
Disk ATLAS Parity OK   Production 131742 Ok  
Disk CMS Degraded Parity   Production 133515 raid5 in a few LUNs, raid6 in the others Disks to be replaced
Disk LHCb Parity Restored Production in March Not Ready 133673 raid6 in all LUNs Started the copy of data from old system to new one, ETA middle~end of this week. Tapes should be ready to stage when when copy is done
Computing farm -   Ready   CMS: configure singularity on only CINECA farm (DONE); ATLAS: FS mounting on farm nodes and restart LSF queues.   Now all CEs in unscheduled downtime - WIP, should be back today or tomorrow

  • EGI: nc
  • FNAL: nc
  • IN2P3: Tomorrow morning, ALICE and LHCb production will be fully moved to CentOS7 WNs. Next tuesday March 13th, IN2P3-CC will in scheduled maintenance. CE and SE will be in downtime and back by the end of the day.
  • JINR: NTR
  • KISTI: NTR
  • KIT: NTR
  • NDGF:
    • Switch maintenance tomorrow morning (07-08 CET), might cut a chunk of ATLAS data and the squid server unavailable for a short time.
  • NL-T1: NTR
  • NRC-KI: nc
  • OSG: NTR
  • PIC: NTR
  • RAL: NTR
  • TRIUMF: Maintenance on site core router on March 4, there were two network connectivity outages and each lasted about 12 minutes.

  • CERN computing services: NTR
  • CERN storage services: EOS-CMS was slow during the week end night due to a find command that was performed, currently being investigated
  • CERN databases: CMSONR DB hosts were restarted today to refresh DNS settings
  • GGUS: NTR
  • Monitoring:
    • Final SAM report for Jan 2018 sent to WLCG office.
  • MW Officer: nc
  • Networks: AGLT2 to LHCONE sites (NDGF/CERN) under investigation, narrowed down to US/ESNet segment
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r18 < r17 < r16 < r15 < r14 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r18 - 2018-03-05 - KateDziedziniewicz
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback